x^3-x^2-12x=0


I think the answer is x=0, x=4, x=-3, but I'm not really sure how I got that...

Found 2 solutions by stanbon, jayanthihemkos@yahoo.com:
Answer by stanbon(75887)

x^3-x^2-12x=0
x(x^2-x-12)=0
Factor to get:
x(x-4)(x+3)=0
x=0 or x=4 or x=-3
==========
Cheers,
Stan H.

Answer by jayanthihemkos@yahoo.com(87)

x^3-x^2-12x=0
Take x as common factor
x(x^2-x -12) = 0
x = 0
x^2 -x -12 = 0
Factorize the above trinomial equation by middle term factorization

x^2 -4x +3x -12 =0
x(x-4) +3(x-4) = 0 (taking common factor)
(x+3)(x-4) = 0 (factors of the trinomial equation)
x+3 = 0
x+3-3 = 0-3
x = -3
x-4 = 0
x-4+4 = 0+4
x = 4

Hence x = 0 , -3, 4 is the solution for the above polynomial equation